Qd=17.5-(1/6)P, Qs=(1/3)P-2.
What is the equilibrium price of a ton of blueberries? How many tons will be provided?How much consumer surplus and producer surplus does this market provide? What are the total gains from trade?
Suppose the government decides that for the good of the country, no one should be allowed to sell mushrooms at a price higher than $33. How many tons of blueberries will be traded? What is the new consumer surplus, the new producer surplus and the new total gains from trade?
What was the “waste” caused by this government policy?
1
Expert's answer
2019-10-17T10:06:56-0400

The equilibrium price of a ton of blueberries and quantity are:

Qd = Qs,

17.5 - (1/6)P = (1/3)P - 2

0.5P = 19.5,

P = 39,

Q = 1/3×39 - 2 = 11 tons.

The consumer surplus is:

CS = 0.5×(17.5×6 - 39)×11 = 363.

The producer surplus is:

PS = 0.5×(39 - 6)×11 = 181.5.

If the price ceiling is set at the price of 33, then there will be a shortage of blueberries, the consumer surplus will increase, and the producer surplus will decrease.